a bullrt of mass 20 g leaves  rifle at an initial speed 100m/s and strikes a target at same level with speed 50 m/s the amount of work done by resistance of air is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Harshita
 
by work - energy theorem,
initial KE - final KE = work performed against air resistance
Initial KE = 1/2 * m * vi^2 
Final KE = 1/2 * m * vf^2
Hence change = 1/2 * 20/1000 * (100^2 - 50^2) = 1/100 * 150 * 50 = 75 J
So work done by bulllet against air resistance = 75 J
Or work performed by air to reduce the energy of bullet = 75 J
 
Regards,
Arun (AskIITians Forum Expert)